# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ Abstract
Distributed stochastic gradient descent (SGD) has attracted considerable recent attention due to its potential for scaling computational resources, reducing training time, and helping protect user privacy in machine learning. However, the staggers and limited bandwidth may induce random computational/communication delays, thereby severely hindering the learning process. Therefore, how to accelerate asynchronous SGD by efficiently scheduling multiple workers is an important issue. In this paper, a unified framework is presented to analyze and optimize the convergence of asynchronous SGD based on stochastic delay differential equations (SDDEs) and the Poisson approximation of aggregated gradient arrivals. In particular, we present the run time and staleness of distributed SGD without a memorylessness assumption on the computation times. Given the learning rate, we reveal the relevant SDDE’s damping coefficient and its delay statistics, as functions of the number of activated clients, staleness threshold, the eigenvalues of the Hessian matrix of the objective function, and the overall computational/communication delay. The formulated SDDE allows us to present both the distributed SGD’s convergence condition and speed by calculating its characteristic roots, thereby optimizing the scheduling policies for asynchronous/event-triggered SGD. It is interestingly shown that increasing the number of activated workers does not necessarily accelerate distributed SGD due to staleness. Moreover, a small degree of staleness does not necessarily slow down the convergence, while a large degree of staleness will result in the divergence of distributed SGD. Numerical results demonstrate the potential of our SDDE framework, even in complex learning tasks with non-convex objective functions.
@@
Distributed optimization, machine learning, stochastic gradient descent, gradient staleness, computational delay, communication bandwidth, stochastic delay differential equations, asynchronous algorithm, event-based signal processing, performance analysis
@@ I Introduction
Deep neural networks (DNNs) have found widespread applications in diverse fields, including computer vision, natural language processing, and emerging areas like textual data analysis [1], in which high-performance multi-GPU computing is enabled by GPU-oriented interconnects such as NVLink[2]. Computer networks, supported by communication protocols [3, 4], often serves as the interconnecting infrastructure enabling efficient resource sharing among network nodes [5], which provides the essential framework for distributed AI systems to communicate and access resources. Stochastic gradient descent (SGD) serves as the fundamental framework for most existing deep learning algorithms, enabling models to learn from extensive databases and attain exceptional performance.
@@
Although parallelizing training can accelerate the training process, the presence of straggling workers can lead to idle time, reducing the speedup potential. In traditional synchronous learning, faster devices are forced to wait for slower ones due to the heterogeneity of the devices. To tackle this issue, asynchronous learning has been proposed, where each worker pushes gradients to the server in an event-triggered manner [7, 8, 9, 10, 11, 12, 13]. As an important advantage of asynchronous SGD, it has been shown to achieve better performance compared to synchronous SGD with respect to wall-clock time [14]. Though asynchronous SGD eliminates waiting overhead, it suffers from higher convergence errors due to gradient staleness, which is induced by the computational delay and communication delay [15, 17]. As a result, stale gradients are usually penalized using a staleness-dependent learning rate in staleness-aware asynchronous SGD [19, 18, 21, 20]. Also, the communication load can be reduced in distributed learning via event-triggered communication, referred to as event-triggered stochastic gradient descent [22, 23, 24, 25, 26].
@@
For the theoretical convergence analysis of SGD, classical bound-type results include [27] and [28]. Existing bound-type results for SGD with stale-gradient include [29, 16, 30, 14, 33, 32]. The key obstacle to deal with in asynchronous learning is gradient staleness. In [29], asynchronous SGD is shown to suffer an asymptotic penalty in convergence rate due to gradient staleness. In the presence of gradient delays, Agarwal and Duchi demonstrate that asynchronous learning can achieve order-optimal convergence results in [30]. As is also presented in [32], asynchronous SGD achieves similar convergence property as the centralized counterpart with bounded communication delay. To tackle this issue of gradient staleness, by carefully tuning the algorithm’s step size, asynchronous learning is shown to still converge the critical set in [34]. In addition to the classical bound type analysis, performance analysis based on continuous approximation by stochastic differential equation also attracted considerable attention in stochastic approximation literature [35, 36]. In [38], the author uses tools from stochastic calculus and asymptotic analysis to provide a precise dynamical description of SGD and its variants, based on which adaptive learning rate policies are presented. In [37], the authors approximated SGD in terms of a multivariate Ornstein-Uhlenbeck process, where precise, albeit only distributional, descriptions of the SGD dynamics are presented in complement to the classical bound type converge analysis. In addition, the authors conduct theoretical analysis on the convergence rates through the continuous approximation by stochastic differential delay equations in [33].
@@
To provide non-asymptotic performance analysis and gain greater insights into the impact of gradient staleness on the convergence of distributed SGD, in this paper, a stochastic approximation-based approach is presented, in complement to the conventional bound type analysis. Similar to [37, 38, 10], we first focus on quadratic objective functions and then extend our discussion to general cases. Further, we aim to derive criteria for the parameter setup of learning rates, the total number of workers, and communication protocol in asynchronous distributed learning with gradient staleness. In the literature, closely related works are [14, 15] in terms of wall-clock run-time analysis, [37, 38] in terms of stochastic differential equation approximation, and [33] in terms of the connection between asynchronous SGD and stochastic delay differential equations (SDDEs). In this work, we aim to extend the run-time analysis for the variants of SGD with exponential processing times [14] to non-memoryless cases, and stochastic differential equation approximation for SGD in [37, 38] to SDDE approximation for asynchronous SGD with gradient staleness. In this work, in complement to the conventional bound-type analysis, we aim to present an SDDE-based approach to reveal the convergence rate of asynchronous SGD. In addition, the run-time and step staleness analysis is further presented. Specifically, the contribution of this paper can be summarized as follows:
@@
  • We present a unified framework to analyze and optimize the convergence process of asynchronous SGD with gradient noise and staleness based on SDDEs. To deal with the non-exponentially distributed computation time, a Poisson approximation for the superposition of independent renewal processes is adopted, based on which we present the probability distribution of the update interval and the step staleness or discrete-time gradient staleness of asynchronous SGD. Based on the runtime and staleness analysis, we bridge the behavior of the asynchronous SGD algorithm and the solutions of the SDDE. In particular, the convergence time or rate of convergence of the SGD algorithm can be determined based on the first hitting time of the stochastic process or characteristic roots of the SDDE. It is interestingly shown that, regardless of the specific distribution of computation time, the expectation of the step staleness in asynchronous SGD without gradient dropout is only determined by the number of workers and the group size.
@@
  • The performance of the SGD algorithm with stale gradient is demonstrated to be closely related to the product of the step staleness, learning rate, and the 2-norm of the Hessian matrix. Our analysis indicates that a small degree of staleness could slightly accelerate the SGD algorithm, while a large degree staleness could be harmful and result in the divergence of asynchronous SGD. Our analyses align with the observations made by existing literature and complement classical bound-type convergence analyses for SGD with stale gradients. Also, it is shown that the presence of the gradient noise contributes to an increase in the average first hitting time. Under certain circumstances, the increase is proven to be proportional to the standard deviation of the Gaussian noise and the square root of the logarithm of the optimization variable’s dimension.
@@
  • An excessive number of activated workers does not necessarily accelerate asynchronous SGD. This is due to the fact that a small number of workers provides fewer gradients while a large number of workers leads to a greater degree of staleness. With a staleness-aware learning rate, the positive effect of the increased number of gradients as the number of workers increases can be offset by the staleness-aware step size. The performance of asynchronous SGD with an excess of workers can be improved by selecting an appropriate group size or discarding outdated gradients. The limited bandwidth leads to an increased gradient staleness or a smaller update frequency depending on the client starting the computation of a new gradient right after finishing the computation or the transmission of the old one. In this case, an excessive number of workers can lead to a large communication delay due to network congestion, resulting in a degradation in the performance of SGD. Further, high-resolution quantization of the gradients with large gradient noise can lead to an overload of network bandwidth, thereby being inefficient due to the waste of communication resources.
  • A uniformly distributed gradient staleness arises in event-triggered SGD. To characterize the convergence rate of event-triggered SGD, the characteristic roots of the SDDE with uniformly distributed staleness are revealed. Similar to asynchronous SGD, it is shown that a small triggering threshold in event-triggered SGD could slightly accelerate event-triggered SGD and reduce the communication load, while a large triggering threshold may lead to a large degree of staleness and therefore lead to the divergence of event-triggered SGD.
@@
$\hat{\tau}$ Gradient delay or continuous-time gradient staleness $\boldsymbol{w}$ Gaussian noise
@@
II System Model
Figure 1: System model.
In this paper, we are interested in centralized distributed learning involving gradient noise and staleness. Specifically, there is a single central parameter server connected with $K$ parallel clients, or workers, via a communication medium as illustrated in Fig. 1. In the conventional synchronous setting, the global model evolves based on gradients from all the $K$ workers at each iteration. To mitigate stragglers in synchronous SGD, asynchronous SGD has been introduced. In the asynchronous setup, each worker can operate independently, fetching the global model and updating its computed gradient. Furthermore, to alleviate the communication overload in distributed learning, event-triggered SGD is presented where the worker updates its gradient to the central server until the difference between the previous gradient and the current one exceeds a given triggering threshold. We introduce the implementation details of asynchronous SGD and event-triggered SGD in Subsections II-A and II-B, respectively, and provide corresponding stochastic approximations.
@@ II-A Asynchronous Distributed Optimization
To mitigate stragglers in synchronous distributed learning, asynchronous distributed learning has emerged as an alternative approach. In the asynchronous setting, each worker operates independently of others, fetching the global model and updating its computed gradient asynchronously. The process unfolds as follows: Firstly, a worker first fetches from the central server the most up-to-date parameters of the model to process the current mini-batch. Subsequently, it computes gradients of the loss function with respect to these parameters. These gradients are then sent to the central server, which then updates the model accordingly. Simultaneously, upon completing the computation, it fetches the updated global model right away, based on which it computes new gradients. [^1] In this paper, we assume that the downlink communication cost is negligible. Also, the time it takes for the central server to update the global model is also assumed to be negligible since the update is based on computed gradients. In this way, each worker could fetch the updated global model and start the computation of a new gradient once it finishes the computation of the old one. In the asynchronous setup, in the asynchronous setup, the gradient utilized to update the global model is often computed based on outdated parameters, leading to gradient staleness, which can impede the convergence of asynchronous SGD. To address this challenge, we explore the following two methods aimed at reducing gradient staleness in asynchronous SGD.
@@
First, in $B$-ASGD, the central server updates the global model once $B$ gradients are collected since the last update, while the workers who fail to update the gradient in the current iteration can continue their computation. In this context, $B$ is referred to as the group size. In particular, the $B$ clients, based on whose gradients the global model updates, start the computation of a new gradient based on the updated global model, to reduce the gradient staleness. Note that the $B$ clients could also start their computation of new gradients right after their gradients are received by the central server. In this case, the asynchronous SGD algorithm is termed $B$-batch-ASGD. In this paper, we generally focus on $B$-ASGD unless otherwise stated. Note that asynchronous SGD reduces to synchronous SGD if $B=K$.
@@
Second, the workers could drop the current computing gradient to avoid a large degree of gradient staleness. In particular, let $\hat{\tau}_{\text{th}}$ denote the continuous-time gradient staleness threshold. More specifically, during the computation of a gradient, if the time since a worker last fetched the global model exceeds $\hat{\tau}_{th}$, the worker stops the computation of the current gradient, fetches the current model, and start the computation of a new gradient. In fact, the worker can also drop outdated gradients according to discrete-time staleness or step staleness. In particular, if the number of updates taken by the central server exceeds a pre-given threshold $\breve{\tau}_{\text{th}}$ since the client last fetches the global model, the client halts the current training and starts the computation of the new gradient based on the current global model. By this means, asynchronous SGD reduces the gradient delay at the expense of a smaller update frequency. In this context, we call it pure asynchronous SGD when $B=1$ and $\hat{\tau}_{\text{th}}=\infty$ or $\breve{\tau}_{\text{th}}=\infty$, also referred to as $1$-ASGD without gradient dropout.
@@
$\!\sum_{\kappa\in\mathcal{K}(t)}\!\eta\!\left[\nabla f\left(\hat{\boldsymbol{x}}({t-\hat{\tau}^{\kappa}(t)}\right))\!+\!\boldsymbol{w}^{\kappa}(t)\right]\!\times\!\mathds{1}\left\{\|\mathcal{K}(t)\|\!=\!B\right\},$
in which $\eta$ is the learning rate, $f(\cdot)$ is the objective function, $\nabla f(\cdot)$ is the gradient function of the objective function, $\boldsymbol{w}^{\kappa}(\cdot)$ is the gradient noise corresponds to Worker $\kappa$’s gradient, and $\hat{\tau}(\cdot)$ is the gradient delay or the continuous-time gradient staleness. More specifically, $\eta$ is a constant with a fixed learning rate or a function of the gradient delay or the step staleness with a staleness-aware learning rate. In addition, $\hat{\tau}(\cdot)$ refers to the time between a worker fetches the global model and updates the computed gradient to the central server and $\hat{\tau}^{\kappa}(t)$ denotes the gradient delay of Worker $\kappa$’s gradient that is delivered to the central server at time $t$. In practice, the gradient delay is joint determined by the computation/communication resources, the group size $B$, and the continuous-time or discrete-time staleness threshold $\hat{\tau}_{th}$ or $\breve{\tau}_{th}$. Also, $\boldsymbol{w}^{\kappa}(\cdot)$ is determined by the mini-batch size of each client. Here, we adopt the general assumption of the Gaussianity of the gradient noise, which arises from the central limit theorem [41]. Also, we assume that the Gaussian distribution is isotropic and the variances of the gradient noise across each worker are identical.
@@
in which $\breve{\tau}_{B}^{\kappa}[j]$ is the iterations taken by the global model between Worker $\kappa$ fetches the global model and finishes the computation of the gradient that contributes to the $j$th update of the global model. In $B$-ASGD, the relationship between the step staleness or discrete-time staleness $\breve{\tau}^{\kappa}[j]$ and $\breve{\tau}_{B}^{\kappa}[j]$ is given by $\breve{\tau}^{\kappa}[j]=\breve{\tau}_{B}^{\kappa}[j]\times B$, since the global model is updated based on $B$ gradients at each iteration. The relationship between the discrete-time gradient staleness $\breve{\tau}^{\kappa}[j]$ and the continuous-time staleness $\hat{\tau}^{\kappa}(t)$ in $B$-ASGD are given by
@@
Let the step staleness of the arrived gradients be denoted by the random variable $\breve{\tau}$. Intuitively, a larger $\mathbb{E}\{\breve{\tau}\}$ corresponds to an increased training error, as discussed in [18]. Also, it is noteworthy that a larger gradient delay does not necessarily result in a greater step staleness.
Let $\boldsymbol{x}(\cdot)$ denote the continuous approximation of $\hat{\boldsymbol{x}}(\cdot)$ with respect to the stochastic differential equation. In this way, the continuous approximation for asynchronous SGD is presented via the following stochastic delay differential equation:

### v2

@@ Abstract
Distributed stochastic gradient descent (SGD) has attracted considerable recent attention due to its potential for scaling computational resources, reducing training time, and helping protect user privacy in machine learning. However, stragglers and limited bandwidth may induce random computational/communication delays, thereby severely hindering the learning process. Therefore, how to accelerate asynchronous SGD (ASGD) by efficiently scheduling multiple workers is an important issue. In this paper, a unified framework is presented to analyze and optimize the convergence of ASGD based on stochastic delay differential equations (SDDEs) and the Poisson approximation of aggregated gradient arrivals. In particular, we present the run time and staleness of distributed SGD without a memorylessness assumption on the computation times. Given the learning rate, we reveal the relevant SDDE’s damping coefficient and its delay statistics, as functions of the number of activated clients, staleness threshold, the eigenvalues of the Hessian matrix of the objective function, and the overall computational/communication delay. The formulated SDDE allows us to present both the distributed SGD’s convergence condition and speed by calculating its characteristic roots, thereby optimizing the scheduling policies for asynchronous/event-triggered SGD. It is interestingly shown that increasing the number of activated workers does not necessarily accelerate distributed SGD due to staleness. Moreover, a small degree of staleness does not necessarily slow down the convergence, while a large degree of staleness will result in the divergence of distributed SGD. Numerical results demonstrate the potential of our SDDE framework, even in complex learning tasks with non-convex objective functions.
@@
Distributed optimization, machine learning, stochastic gradient descent, gradient staleness, stochastic delay differential equations, asynchronous distributed optimization, straggler problem.
@@ I Introduction
Deep neural networks (DNNs) have found widespread applications in diverse fields, including computer vision, natural language processing, and emerging areas like textual data analysis [1], in which high-performance multi-GPU computing is enabled by GPU-oriented interconnects such as NVLink[2]. Computer networks, supported by communication protocols [3, 4], often serve as the interconnecting infrastructure enabling efficient resource sharing among network nodes [5], which provides the essential framework for distributed AI systems to communicate and access resources. Stochastic gradient descent (SGD) serves as the fundamental framework for most of the existing deep learning algorithms, enabling models to learn from extensive databases and attain exceptional performance.
@@
Parallelizing training can accelerate the overall training process. However, due to the need for synchronized updates, the completion time of each iteration is constrained by the slowest, so-called “straggling” workers [6, 7]. This can negatively impact the convergence of the algorithm, as the presence of these stragglers leads to idle time, thereby reducing the potential speedup. In traditional synchronous learning, faster devices are forced to wait for slower ones due to the heterogeneity of the devices. To tackle this issue, asynchronous learning has been proposed, where each worker pushes gradients to the server in an event-triggered manner [9, 10, 11, 12, 13, 14, 15]. As an important advantage of asynchronous SGD (ASGD), it has been shown to achieve better performance compared to synchronous SGD with respect to wall-clock time [16]. Though ASGD eliminates waiting overhead, it suffers from higher convergence errors due to gradient staleness, which is induced by the computational delay and communication delay [17, 18]. As a result, stale gradients are usually penalized using a staleness-dependent learning rate in staleness-aware ASGD [20, 19, 22, 21]. Also, communication load can be reduced in distributed learning via event-triggered communication, referred to as event-triggered stochastic gradient descent [23, 24, 25, 26, 27].
@@
For the theoretical convergence analysis of SGD, classical bound-type results include [28] and [29]. Existing bound-type results for SGD with stale-gradient include [30, 31, 32, 16, 34, 33, 37, 38, 36, 35, 39]. The key obstacle to deal with in asynchronous learning is gradient staleness. In [30], ASGD is shown to suffer an asymptotic penalty in convergence rate due to gradient staleness. In the presence of gradient delays, Agarwal and Duchi demonstrate that asynchronous learning can achieve order-optimal convergence results in [32]. As is also presented in [33], ASGD achieves similar convergence property as the centralized counterpart with bounded communication delay. To tackle this issue of gradient staleness, by carefully tuning the algorithm’s step size, asynchronous learning is shown to still converge the critical set in [40]. In addition to the classical bound type analysis, performance analysis based on continuous approximation by stochastic differential equation also attracted considerable attention in stochastic approximation literature [41, 42]. In [43], the author uses tools from stochastic calculus and asymptotic analysis to provide a precise dynamical description of SGD and its variants, based on which adaptive learning rate policies are presented. In [44], the authors approximated SGD in terms of a multivariate Ornstein-Uhlenbeck (OU) process, where precise, albeit only distributional, descriptions of the SGD dynamics are presented in complement to the classical bound type converge analysis. In addition, the authors conduct theoretical analysis on the convergence rates through the continuous approximation by stochastic differential delay equations in [34].
@@
Motivated by the straggler problem, and to provide non-asymptotic performance analysis and gain greater insights into the impact of gradient staleness on the convergence of distributed SGD, in this paper, a stochastic approximation-based approach is presented, in complement to the conventional bound type analysis. Similar to [44, 43, 12], we first focus on quadratic objective functions and then extend our discussion to general cases. Further, we aim to derive criteria for the parameter setup of learning rates, the total number of workers, and communication protocol in asynchronous distributed learning with gradient staleness. In the literature, closely related works are [16, 17] in terms of wall-clock run-time analysis, [44, 43] in terms of stochastic differential equation approximation, and [34] in terms of the connection between ASGD and stochastic delay differential equations (SDDEs). In this work, we extend the run-time analysis of SGD variants with exponential processing times [16] to non-memoryless cases. We further extend the stochastic differential equation (SDE) approximation of SGD, developed in [44, 43], to an SDDE approximation that accounts for gradient staleness in ASGD. Furthermore, inspired by the stationary distribution analysis of SGD in [44], we perform a first hitting time analysis for SGD and a stationary distribution analysis for ASGD. In addition to [34], we bridge the SDDE solution to ASGD dynamics with respect to the wall-clock time, and we also provide a detailed analysis of the characteristic roots of the SDDE, highlighting how factors such as the learning rate, the number of workers, and gradient staleness influence the ASGD convergence rate.
Figure 1: System model.
In this paper, we develop an SDDE-based approach, through which we bridge ASGD dynamics to the solution of an SDDE. Specifically, by presenting the characteristic root of the SDDE, we provide an insightful argument on the relationship between the convergence rate and gradient staleness, in complement to existing bound type analysis. Moving away from the memorylessness assumptions on computational times, the run-time and step staleness analysis in SDDE is further presented. This approach enables us to reveal how factors such as learning rate, gradient staleness, number of workers, and communication protocol design influence ASGD performance, as well as presenting protocol design criteria. Specifically, the contribution of this paper can be summarized as follows:
@@
  • We present a unified framework to analyze and optimize the convergence process of ASGD with gradient noise and staleness based on SDDEs. To deal with the non-exponentially distributed computation time, a Poisson approximation for the superposition of independent renewal processes is adopted, based on which we present the probability distribution of the update interval and the step staleness or discrete-time gradient staleness of ASGD. Based on the runtime and staleness analysis, we bridge the behavior of the ASGD algorithm and the solutions of the SDDE. In particular, the convergence time or rate of convergence of the SGD algorithm can be determined based on the first hitting time of the stochastic process or characteristic roots of the SDDE.
@@
  • The performance of the SGD algorithm with stale gradient is demonstrated to be closely related to the product of the step staleness, learning rate, and the 2-norm of the Hessian matrix. Our analysis indicates that a small degree of staleness could slightly accelerate the SGD algorithm, while a large degree of staleness could be harmful and result in the divergence of ASGD. Our analyses align with the observations made by existing literature and complement classical bound-type convergence analyses for SGD with stale gradients. It is further shown that gradient staleness slows down the escape of the ASGD algorithm from saddle points. We further provide a theoretical analysis of the impact of uniformly distributed gradient staleness on the convergence rate of ASGD.
  • It is further revealed that reducing gradient staleness comes at the cost of a lower gradient update rate. It is interestingly shown that, regardless of the specific distribution of computation time, the expected step staleness in ASGD without gradient dropout is only determined by the number of workers and the group size. Also, it is shown that the presence of the gradient noise contributes to an increase in the average first hitting time. Under certain circumstances, the increase is proven to be proportional to the standard deviation of the Gaussian noise and the square root of the logarithm of the optimization variable’s dimension. This work also shows that ASGD has a higher error floor compared to synchronous SGD. However, this issue can be alleviated by decreasing the level of asynchrony as ASGD nears the local minimum.
@@
  • In practice, the optimal choice of worker number can be determined according to the learning rate and the $2$-norm of the Hessian matrix of the objective function. An excessive number of activated workers does not necessarily accelerate ASGD. This is due to the fact that a small number of workers provides fewer gradients while a large number of workers leads to a greater degree of staleness. The performance of ASGD with an excess of workers can be improved by selecting an appropriate group size or discarding outdated gradients. The limited bandwidth leads to an increased gradient staleness or a smaller update frequency depending on whether the client starts computing a new gradient immediately after completing or transmitting the previous one. In this case, an excessive number of workers can lead to a large communication delay due to network congestion, resulting in a degradation in the performance of SGD. Further, high-resolution quantization of the gradients with large gradient noise can lead to an overload of network bandwidth, thereby being inefficient due to the waste of communication resources.
  • ASGD offers an effective solution to address the straggler problem arising in synchronous SGD. While ASGD may be hindered by stale gradients, this paper demonstrates that, through the careful selection of the number of workers or by designing appropriate protocols, step staleness can be effectively managed. Also, it is possible to leverage stale gradients to accelerate the convergence of ASGD. Additionally, ASGD benefits from a higher gradient arrival rate compared to synchronous SGD, which can lead to better convergence rates in terms of wall-clock time.
II System Model
@@
$\hat{\tau}$ Gradient delay or continuous-time gradient staleness $\boldsymbol{w}$ The gradient noise
@@
In this paper, we are interested in centralized distributed learning involving gradient noise and staleness. Specifically, there is a single central parameter server connected with $K$ parallel clients, or equivalently workers, via a communication medium as illustrated in Fig. 1. In the conventional synchronous setting, the global model evolves based on gradients from all the $K$ workers at each iteration. To mitigate stragglers in synchronous SGD, ASGD has been introduced. In the asynchronous setup, each worker can operate independently, fetching the global model and updating its computed gradient. Furthermore, to alleviate the communication overload in distributed learning, event-triggered SGD is presented where the worker pushes its gradient to the central server until the difference between the previous gradient and the current one exceeds a given triggering threshold. We introduce the implementation details of ASGD in Subsection II-A, and provide a corresponding stochastic approximation. In Subsection II-B, key mathematical concepts used in this paper are presented.
@@ II-A Asynchronous Distributed Optimization
Let us first introduce the objective function in distributed SGD. Let $\boldsymbol{x}\in\mathbb{R}^{N}$ be the parameters of the model to be trained. Let $f_{i}(\boldsymbol{x})$ be the loss incurred by $\boldsymbol{x}$ at the $i^{\text{th }}$ data point $\xi_{i}\sim\mathcal{D}$, in which $\mathcal{D}$ is the set of the training data. The objective is to minimize the generalization error
$f(\boldsymbol{x})=\mathbb{E}_{\xi_{i}\sim\mathcal{D}}\left\{f_{i}(\boldsymbol{x})\right\},$ (1)
@@
where $\mathbb{E}\{\cdot\}$ denotes expectation. To mitigate stragglers in conventional synchronous distributed learning, asynchronous distributed learning has emerged as an alternative approach. In the asynchronous setting, each worker operates independently of others, fetching the global model and updating its computed gradient asynchronously. The process unfolds as follows: Firstly, a worker fetches from the central server the most up-to-date parameters of the model to process the current mini-batch. Subsequently, it computes gradients of the loss function with respect to these parameters. These gradients are then sent to the central server, which then updates the model accordingly. Simultaneously, upon completing the computation, it fetches the updated global model right away, based on which it computes new gradients. [^1] In this paper, we assume that the downlink communication cost is negligible. Also, the time it takes for the central server to update the global model is also assumed to be negligible since the update is based on computed gradients. In this way, each worker could fetch the updated global model and start the computation of a new gradient once it finishes the computation of the old one. In the asynchronous setup, the gradient utilized to update the global model is often computed based on outdated parameters, leading to gradient staleness, which can impede the convergence of ASGD. To address this challenge, we explore the following two methods aimed at reducing gradient staleness in ASGD.
@@
First, in $B$-ASGD, the central server updates the global model once $B$ gradients are collected since the last update, while the workers who fail to push the gradient in the current iteration can continue their computation. In this context, $B$ is referred to as the group size. In particular, the $B$ clients, based on whose gradients the global model updates, start the computation of a new gradient based on the updated global model, to reduce the gradient staleness. Note that the $B$ clients could also start their computation of new gradients right after their gradients are received by the central server. In this case, the ASGD algorithm is termed $B$-batch-ASGD. In this paper, we generally focus on $B$-ASGD unless otherwise stated. Note that ASGD reduces to synchronous SGD if $B=K$ [^2] In this paper, B-ASGD follows the same algorithm as K-async SGD described in [16], with additional features such as gradient dropout to manage gradient staleness. Further, as our study uses different analytical methods and assumptions, we use distinct terminology to differentiate the two approaches.
@@
Second, the workers could drop the current computing gradient to avoid a large degree of gradient staleness. In particular, let $\hat{\tau}_{\text{th}}$ denote the continuous-time gradient staleness threshold. More specifically, during the computation of a gradient, if the time since a worker last fetched the global model exceeds $\hat{\tau}_{th}$, the worker stops the computation of the current gradient, fetches the current model, and start the computation of a new gradient. In fact, the worker can also drop outdated gradients according to discrete-time staleness or step staleness. In particular, if the number of updates taken by the central server exceeds a pre-given threshold $\breve{\tau}_{\text{th}}$ since the client last fetched the global model, the client halts the current training and starts the computation of the new gradient based on the current global model. By this means, ASGD reduces the gradient delay at the expense of a smaller gradient update frequency. In this context, we call it pure ASGD when $B=1$ and $\hat{\tau}_{\text{th}}=\infty$ or $\breve{\tau}_{\text{th}}=\infty$.
@@
$\sum_{\kappa\in\mathcal{K}(t)}\eta\Big{[}\nabla f\left(\hat{\boldsymbol{x}}({t-\hat{\tau}^{\kappa}(t)}\right))+\boldsymbol{w}^{\kappa}(t)\Big{]}\times\mathds{1}\left\{\|\mathcal{K}(t)\|=B\right\},$
in which $\eta$ is the learning rate, $f(\cdot)$ is the objective function, $\nabla f(\cdot)$ is the gradient function of the objective function, $\boldsymbol{w}^{\kappa}(\cdot)$ is the gradient noise corresponds to Worker $\kappa$’s gradient, and $\hat{\tau}(\cdot)$ is the gradient delay or the continuous-time gradient staleness. More specifically, $\eta$ is a constant with a fixed learning rate. With a staleness-aware learning rate, $\eta$ can be a function of the gradient delay, or equivalently, continuous-time gradient staleness. Alternatively, $\eta$ can also be a function of the step staleness, or equivalently, discrete-time gradient staleness.. In addition, $\hat{\tau}(\cdot)$ refers to the time between a worker fetches the global model and pushes the computed gradient to the central server, and $\hat{\tau}^{\kappa}(t)$ denotes the gradient delay of Worker $\kappa$’s gradient that is delivered to the central server at time $t$. In practice, the gradient delay is jointly determined by the computation/communication resources, the group size $B$, and the continuous-time or discrete-time staleness threshold $\hat{\tau}_{th}$ or $\breve{\tau}_{th}$. Also, $\boldsymbol{w}^{\kappa}(\cdot)$ is determined by the mini-batch size of each client. Here, we adopt the general assumption of the Gaussianity of the gradient noise, which arises from the central limit theorem [45]. Also, we assume that the Gaussian distribution is isotropic and the variances of the gradient noise across each worker are identical. Next, we present the following two assumptions on the objective function and the gradient noise.
Assumption 1
The objective function $f(x)$ is assumed to be twice-differentiable at $\boldsymbol{x}\in\mathbb{R}^{N}$, in which $\boldsymbol{x}$ represents the parameters of the model to be trained as defined in Eq. (1).
Assumption 2
The gradient noise $\boldsymbol{w}$ is assumed to follow an isotropic Gaussian distribution.
Under Assumption 1, we approximate the twice-differentiable objective function $f(\boldsymbol{x})$ using Taylor expansion. More precisely, we have $f(\boldsymbol{x})=f(\boldsymbol{x}_{0})+(\boldsymbol{x}-\boldsymbol{x}_{0})^{T}\nabla f(\boldsymbol{x}_{0})+(\boldsymbol{x}-\boldsymbol{x}_{0})^{T}\boldsymbol{H}(\boldsymbol{x}_{0})(\boldsymbol{x}-\boldsymbol{x}_{0})+\mathcal{O}(\delta^{3}),$ by Taylor expansion under Assumption 1 for $\|\boldsymbol{x}-\boldsymbol{x}_{0}\|<\delta$ with some small positive $\delta$ [46, Chapter 3.3]. The assumption of twice differentiability is made to facilitate subsequent analysis, particularly for discussing the impact of staleness on convergence, based on the positive definiteness of the Hessian matrix.
@@
in which $\breve{\tau}_{B}^{\kappa}[j]$ is the iterations taken by the global model between Worker $\kappa$ fetches the global model and finishes the computation of the gradient that contributes to the $j$th update of the global model. In $B$-ASGD, the relationship between the step staleness or discrete-time staleness $\breve{\tau}^{\kappa}[j]$ and $\breve{\tau}_{B}^{\kappa}[j]$ is given by $\breve{\tau}^{\kappa}[j]=\breve{\tau}_{B}^{\kappa}[j]\times B$, since the global model is updated based on $B$ gradients at each iteration. The relationship between the discrete-time gradient staleness $\breve{\tau}^{\kappa}[j]$ and the continuous-time staleness $\hat{\tau}^{\kappa}(t)$ in $B$-ASGD is given by
@@
Let the step staleness of the arrived gradients be denoted by the random variable $\breve{\tau}$. Intuitively, a larger $\mathbb{E}\{\breve{\tau}\}$ corresponds to an increased training error, as discussed in [19]. Also, it is noteworthy that a larger gradient delay does not necessarily result in a greater step staleness. Note that continuous-time gradient staleness and gradient delay refer to the same concept: the time interval between when a worker fetches the global model and when the computed gradient is pushed back to the central server [47]. This value is a non-negative real number. In addition, “discrete-time gradient staleness” and “step staleness” refer to the same concept [13]. This term describes the number of iterations taken by the global model between when a worker fetches the model and when it completes the computation of the gradient. It is a non-negative integer.
Let $\boldsymbol{x}(\cdot)$ denote the continuous-approximation of $\hat{\boldsymbol{x}}(\cdot)$ with respect to the stochastic differential equation. In this way, the dynamics of ASGD can be written as an Euler–Maruyama approximation of the following stochastic delay differential equation: